OpenStudy (anonymous):

factor each completely 15x2 - 18y2

OpenStudy (luigi0210):

Uhm, the only thing you can factor is a 3.. \[\LARGE 3(5x^2-6y^2)\]

OpenStudy (jiteshmeghwal9):

\[15x^2-18y^2\]take 3 common\[3(5x^2-6y^2)\]\[a^2-b^2=(a+b)(a-b)\]so,\[(5x^2-6y^2)=(5x+6y)(5x-6y)\]\[3(5x+6y)(5x-6y)\]